# TinyMyo: a Tiny Foundation Model for Flexible EMG Signal Processing at the Edge

### v1

@@ Abstract
Surface electromyography (EMG) is a non-invasive sensing modality used in several domains, including biomechanics, rehabilitation, prosthetic control, and emerging human-machine interaction paradigms. Despite decades of use, significant challenges remain in achieving robust generalization across subjects, recording systems, and acquisition protocols. To tackle these challenges, foundation models (FMs) are gaining traction when targeting end-to-end applications based on EMG signals. Yet, existing EMG FMs remain limited to single downstream tasks and lack deployability on embedded platforms.
In this work, we present TinyMyo, a lightweight FM based on a Transformer encoder architecture. The model is pre-trained in a self-supervised manner on publicly available datasets and achieves high reconstruction fidelity with only 3.6M parameters.
With minimal task-specific head adaptations, the same backbone is used to tackle multiple downstream tasks, leveraging datasets acquired from diverse sensing locations and hardware platforms. We demonstrate generalization across hand gesture classification, hand kinematic regression, speech production and recognition, with performance comparable to or surpassing the state of the art (SoA), and model size below 5M parameters. We achieve SoA results compared to previous FM-based works on the NinaPro DB5 ($89.4\pm 0.16\%$), UCI-EMG ($97.56\pm 0.32\%$), and EPN-612 ($96.74\pm 0.09\%$) datasets.
We report, to the best of our knowledge, the first deployment of an EMG FM on an ultra-low-power microcontroller (GAP9), achieving an average power envelope of $36.45\text{mW}$.
By open-sourcing the pre-trained and the downstream task architectures (https://github.com/pulp-bio/BioFoundation), we aim to provide a flexible resource that can accelerate future research and serve as a common foundation for the EMG community.
@@ 1 Introduction
Electromyography (EMG) measures variations in the electrical activity of muscles during contractions, reflecting the motor-neuron activation patterns [farina_2023_rev]. EMG can be acquired using invasive methods (needle EMG), suitable for a detailed characterization of individual motor units, or using non-invasive, surface-based techniques, to capture the global activity of muscle fibres  [farina_2023_rev, avila_2023_rev].
Surface EMG has found extensive application across a wide range of domains, including biomechanical analysis and rehabilitation [avila_2023_rev], human–robot interaction [xiong_emgrobot_2024], prosthetic control [cimolato_2022emglower, marinelli_2023emgupper], and emerging human–computer interaction interfaces [generic_neuromotor_interface_2025]. However, despite decades of widespread use, many important challenges remain.
First, EMG signals are highly affected by noise, motion artifacts, and cross-talk from adjacent muscles [boyer_emg_noise]. Also, acquisition systems present large variability in terms of number and arrangement of electrodes, sensor geometry, sampling rates, amplification stages, and filtering pipelines [merletti_2020_tutorial]. These sources of variability pose challenges to the design of models that can generalize across heterogeneous hardware configurations. In addition, EMG recordings are strongly influenced by both inter-subject and intra-subject variability [xiong_2021_review]. Differences in anatomy, subcutaneous fat, electrode placement, skin impedance, and day-to-day physiological or behavioural factors lead to substantial shifts in the EMG distribution, posing an additional challenge to generalization across users or recording sessions [ni_2024_survey].
Foundation Models (FM) offer a great promise to address these challenges. Inspired by the success of self-supervised pre-training in vision, language, and speech, as well as by emerging applications in biosignals [gu_2025_foundationsurvery], the core hypothesis is that large-scale datasets can be leveraged to learn high-level, platform and subject-agnostic signal representations. For EMG, such models could provide a unified backbone for many downstream tasks, building upon a robust, device-agnostic feature extraction, improving generalization across devices and subjects. However, progress in FMs of EMG has been limited by the scarcity of large sEMG corpora, and only with recent large-scale datasets (collected from thousands of users) [salter_emg2pose_2024] the exploration of generalized FMs of EMG became viable.
In this context, existing EMG FMs present significant limitations. Current models mostly focus on a single task (typically gesture classification) [chen_2025_physiowave, chen_2025_waveformer]. Furthermore, the size of EMG FMs remains large, and no prior work demonstrated the deployment of an EMG FM on embedded hardware. This motivates the need for efficient FM architectures that meet the memory and computational constraints of embedded devices [benatti_2025_emgreview].
We address these limitations by proposing the following contribution.
@@
  • Unlike previous EMG FMs, which are typically designed around a single downstream task or rely on large, domain-agnostic encoders, TinyMyo generalizes across heterogeneous sEMG datasets and sensor configurations. Our architecture follows the general Transformer design adopted in prior EMG works [chen_2025_physiowave, chen_2025_waveformer], formulated in a unified EMG FM able to support a wide range of downstream tasks using the same pretrained backbone. We demonstrate the versatility of the pre-trained encoder through minimal task-specific head adaptations across diverse downstream tasks, including hand-gesture classification on the Ninapro DB5 [ninapro_db5], EPN-612 [benalcazar_emg-epn-612_2020], UC Irvine [emg_data_for_gestures_481] and Meta’s Generic Neuromotor Interface [generic_neuromotor_interface_2025] datasets, hand-kinematic regression on the Ninapro DB8 [ninapro_db8], and speech production and discrimination on the Gaddy Silent Speech Dataset [gaddy-klein-2020-digital]. Our model achieves competitive performances across all downstream tasks, either surpassing SoA or guaranteeing a comparable accuracy with a sub-5M class model size.
  • We demonstrate for the first time the deployment of an EMG FM on an ultra-low-power microcontroller (GAP9), achieving an average power envelope of $36.45\text{mW}$.
  • We open source the model and weights to enable reproducibility and further research by the community.
@@ 2 Related Works
Several efforts have been made to develop FMs for biosignals, and most of the works focused on electrocardiography (ECG), photoplethysmography (PPG), and electroencephalography (EEG) [ni_2024_survey]. In contrast, research dedicated to EMG FMs remains limited.
@@
Turgut et al. [turgut_2024_otis] introduced OTiS, a foundation model with 45 M parameters pre-trained on a broad collection of modalities, including ECG, temperature, audio, electromechanical signals, EEG, and financial time series. Their work demonstrated the feasibility of handling heterogeneous time-series data, including diverse input types, sampling rates, and channel configurations, within a unified pre-training framework. In the original study, EMG is used exclusively for multi-class muscular disease classification. Subsequent works by Cheng et al. [chen_2025_physiowave, chen_2025_waveformer] evaluated the same model on hand-gesture classification using the EPN612 [benalcazar_emg-epn-612_2020], NinaPro DB5 [ninapro_db5], NinaPro DB6 [ninapro_db6], and UCI EMG [emg_data_for_gestures_481] datasets, demonstrating high accuracy across all datasets.
Goswami et al. proposed MOMENT [goswami_2024_moment], a 350-M parameter FM pre-trained on a wide range of domains and temporal resolutions, including electrical signals, traffic, weather, disease progression, motion capture, simulated data, and ECG (among biopotentials). In the original work, EMG is not included in the pre-training corpus nor in the downstream evaluations. However, recent studies by Cheng et al. [chen_2025_waveformer, chen_2025_physiowave] assessed MOMENT on the hand-gesture classification task using the EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG datasets, demonstrating the applicability of MOMENT as an EMG FM.
@@
PhysioWave [chen_2025_physiowave] introduced a modality-specific FM for EMG and ECG. The authors further propose a unified multimodal framework combining these encoders with existing pre-trained EEG models, leveraging a multi-scale wavelet-transformer architecture, using wavelet decomposition to capture time-frequency features at different scales. For the EMG-specific model, the primary downstream task is hand-gesture classification, evaluated on NinaPro DB5, EPN612, and UCI EMG. PhysioWave achieved a performance comparable to Moment and OTIS with a much smaller parameter count (5M for the smaller model).
The same authors proposed WaveFormer [chen_2025_waveformer], a lightweight Transformer model tailored for sEMG-based gesture recognition, with a model size of only 3.1 M parameters. To ensure the extraction of frequency-domain features remains compact, the model integrates a learnable wavelet transform based on multi-level wavelet decomposition and depthwise separable convolutions. Downstream evaluations are conducted on the EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG datasets, further improving the performance over PhysioWave.
@@
More recently, Jiang et al. [jiang_2025_towards] introduced PhysioOmni, a multi-modal FM designed to handle a wide range of physiological signals such as EEG, ECG, EOG and EMG. For each modality, the authors use the same encoder as in LaBraM [jiang_2024_labram] (5.85.8M parameters per modality, considering its smallest-Base version). They adopt a decoupled multimodal tokenizer and masked signal modelling to ensure robustness under arbitrary missing modalities during inference. During pre-training, PhysioOmni relies on a fixed set of modalities (EEG, EOG, and EMG), which makes the architecture unsuitable for pre-training when one modality is missing. For downstream evaluation, the authors include 5-class sleep stage classification on the HMC [alvarez_2021_hmc] dataset and gait kinematic regression on the FBM [brantley_2018_fbm] dataset.
@@
In this context, existing EMG FMs appear limited by the small number of explored downstream tasks, as well as by the lack of deployability on embedded platforms. TinyMyo is the first EMG FM to cover a broader range of downstream tasks with SoA performance, as will be extensively described in Section 4, while also being deployed on an ultra-low-power MCU.
@@ 3 Methods
Figure 1: Overview of the TinyMyo architecture. (a) Pre-training framework: the input signal is first tokenized using a channel-independent patching strategy with random masking, processed by eight bidirectional Transformer Encoder blocks, and reconstructed by a Lightweight Decoder. (b) Downstream architectures: after pre-training, the Decoder is removed and replaced with task-specific heads. For Hand-Gesture Classification and Kinematic Regression, a linear head is used to produce output classes or joint angles. For speech-related tasks, the EMG input is first downsampled through residual convolutional blocks and then passed through the pre-trained Transformer Encoder, followed by a linear head. The output is text for the Speech Recognition task. For the the Speech Production task, a vocoder (HiFi-GAN) is added to produce audio from the predicted MFCC features. (c) Model Deployment: the architecture for the hand-gesture classification task is implemented on the GAP9 microcontroller.
@@ 3.1.1 Tokenization
TinyMyo adopts a Channel-Independent patching strategy, following methodologies effectively applied in time-series forecasting [nietime] and vision [bao2024_channel_visiontransformersimage]. Unlike other recent biosignal (EEG) foundation models that mix spatial information during the tokenization phase via 2D convolutions that span multiple channels (e.g., FEMBA uses a $2\times 16$ kernel [tegon_2025_femba]) or project them into a unified latent space (e.g., LUNA [doner2025luna]), we strictly maintain per-channel granularity. This design choice is grounded in the physiological nature of EMG recordings, where each channel corresponds to a distinct anatomical location. By treating the channels independently at the embedding stage (using a $1\times L$ kernel), the model allows the subsequent transformer layers to learn inter-channel dependencies via attention, a strategy proven effective for biosignal decoding [dimofte_2025_cerebro].
Formally, let the input EMG waveform be denoted as $\mathbf{X}\in\mathbb{R}^{T\times C}$, with $T$ timesteps and $C$ channels. We apply a sliding window with patch length $L$ and stride $S$ independently to each channel. This yields a grid of patches $\mathbf{P}\in\mathbb{R}^{C\times N_{p}\times L}$, where $N_{p}=\lfloor(T-L)/S\rfloor+1$ is the number of temporal patches per channel.
@@
To form the input sequence for the Transformer, we flatten the channel and temporal dimensions, resulting in a total sequence length of $N=C\times N_{p}$. Each patch $\mathbf{P}_{c,i}$ (where $c$ is the channel index and $i$ is the patch index) is mapped to a latent embedding of dimension $d_{e}$ via a shared learnable linear projection $\mathbf{W}_{\mathrm{proj}}\in\mathbb{R}^{d_{e}\times L}$. The resulting token sequence $\mathbf{E}\in\mathbb{R}^{N\times d_{e}}$ consists of embeddings:
$\mathbf{E}_{k}=\mathbf{W}_{\mathrm{proj}}\,\mathbf{P}_{c,i}^{\top},$ (1)
where the token index $k$ corresponds to the flattened arrangement of channel-time pairs $(c,i)$. We do not add learnable positional embeddings; instead, we inject positional information via Rotary Position Embeddings (RoPE) [su_2023_roformerenhancedtransformerrotary] within the attention blocks.
@@ 3.1.2 Architecture
The Encoder (see Fig. 1) consists of 8 identical layers, based on pre-LayerNorm Transformer encoder [vaswani_2023_attentionneed, xiong_2020_layernormalizationtransformerarchitecture], with 3 attention heads and an embedding dimension of 192, analogous to the configuration used in EfficientViT [liu2023efficientvitmemoryefficientvision].
We adopt a RoPE approach, which introduces position information by rotating query and key vectors in 2D planes instead of adding a position vector; such a method has already been adopted for biosignals [wang_2024_eegpt, chen_2025_physiowave, chen_2025_waveformer].
During pre-training, the Encoder’s output is passed to a lightweight Decoder (3.9k parameters), consisting of a linear layer that outputs a sequence $\mathbf{\hat{P}}$, which is a reconstruction of the original patch sequence $\mathbf{P}$. Our approach draws inspiration from SimMIM [xie2022simmimsimpleframeworkmasked], which demonstrates that a minimal reconstruction head (e.g., a single linear layer) is sufficient for effective self-supervised pre-training.
The use of a lightweight decoder is intentional: placing the full reconstruction burden on the encoder forces it to learn semantically rich and task-agnostic representations, rather than relying on a high-capacity decoder to compensate for missing information. This design strategy is supported by recent survey analyses in the masked image modeling (MIM) literature: shallow decoders tend to limit shortcut learning and encourage the formation of stronger latent representations within the encoder [zhang2022surveymaskedautoencoderselfsupervised]. As a result, the learned features exhibit improved generalization across downstream tasks and recording conditions. Further details on the model’s parameters are reported in Table 1 (left).
@@ 3.1.3 Pre-training objectives
We adopt a random-masking approach [xie2022simmimsimpleframeworkmasked], where a random subset $\mathcal{M}$ of tokens is replaced by a learnable $\mathcal{[MASK]}$ token. Applying the masking over each patch enforces contextual inference over tens of milliseconds (physiologically meaningful burst segments), rather than trivial gap filling. We adopt a masking ratio of $50\%$ as in previous works [dimofte_2025_cerebro, tegon_2025_femba].
We employ the Smooth L1 Loss function as reconstruction objective between masked and original patches, as in [chen_2025_physiowave]. We define the following loss components:
@@
Additional details on pre-training set-up are provided in Table 1 (right).
Table 1: Hyperparameters for masked EMG pre-training.
Hyperparameter Value
Transformer Encoder
Timesteps 1000
Patch size $(H{\times}W)$ $1\times 20$
Channels 16
Embed dim 192
Layers 8
Heads 3
QKV bias Yes
QK norm No
MLP ratio 4.0
MLP size 768
Attn drop 0.1
Proj drop 0.1
Drop-path 0.1
Decoder
Decoder embed dim 192
Setup Value
Batch size 512
Peak LR $1\!\times\!10^{-4}$
Minimum LR $1\!\times\!10^{-6}$
Optimizer $(\beta_{1},\beta_{2})$ AdamW (0.9,0.98)
LR schedule Cosine
Weight decay 0.01
Epochs / warm-up 50 / 10
Grad accum batches 8
Grad clip 3
Mask ratio 0.5
Max seq length 1000
LR: learning rate
All experiments are implemented in Python 3.10 using PyTorch Lightning and Hydra for modularity, configurability, and reproducibility. The proposed foundation model is trained on the CSCS Alps HPC infrastructure using NVIDIA GH200 GPUs, employing a single node with 4x NVIDIA GH200 GPUs in Distributed Data Parallel (DDP) mode.

### v2

@@ Abstract
Objective: Surface electromyography (EMG) is a non-invasive sensing modality widely used in biomechanics, rehabilitation, prosthetic control, and human-machine interfaces. Despite decades of use, achieving robust generalization across subjects, recording systems, and acquisition protocols remains challenging. While foundation models (FMs) are gaining traction for EMG, existing approaches remain limited to single downstream tasks and lack deployability on embedded platforms. This work addresses these limitations. Methods: We present TinyMyo, a lightweight FM based on a Transformer encoder architecture. The model is pre-trained in a self-supervised manner using masked reconstruction on publicly available datasets. With only 3.6M parameters, TinyMyo is designed to support multiple downstream tasks through minimal task-specific head adaptations. Results: We demonstrate generalization across hand gesture classification, hand kinematic regression, speech production and speech recognition, with performance comparable to or surpassing the state of the art (SoA), and model size below 5M parameters. We achieve SoA results compared to previous FM-based works on the NinaPro DB5 (89.4%), UCI-EMG (97.56%), and EPN-612 (96.74%) datasets. We demostrate the first-time deployment of an EMG FM on an ultra-low power microcontroller (GAP9), with an inference time of 0.785 s, energy of 44.91 mJ and power envelope of 57.18 mW. Conclusion: TinyMyo demonstrates that compact, self-supervised EMG FM can guarantee strong generalization across multiple downstream tasks while remaining compatible with low-power edge devices. Significance: TinyMyo is the first EMG FM for ultra-low-power edge-devices, enabling scalable and energy-efficient sensing for motor intent decoding, neuromuscular assessment, and biosignal-driven human–machine interaction.
@@ 1 Introduction
Electromyography (EMG) measures variations in the electrical activity of muscles during contractions, reflecting the motor-neuron activation patterns [15]. EMG can be acquired using invasive methods (needle EMG), suitable for a detailed characterization of individual motor units, or using non-invasive, surface-based techniques, to capture the global activity of muscle fibres  [15, 3].
Surface EMG has found extensive application across a wide range of domains, including biomechanical analysis and rehabilitation [3], human–robot interaction [55], prosthetic control [35], and emerging human–computer interaction interfaces [26]. However, despite decades of widespread use, many important challenges remain.
EMG signals are strongly affected by noise, motion artifacts, cross-talk, and acquisition variability across electrode layouts, hardware platforms, and acquisition pipelines [8, 36]. In addition, both inter-subject and intra-subject factors [56], including anatomy, electrode placement, skin impedance, and physiological variability lead to substantial distribution shifts, hindering generalization [37].
Foundation Models (FM) offer a great promise to address these challenges. Inspired by the success of self-supervised pre-training in vision, language, and speech, as well as by emerging applications in biosignals [23], the core hypothesis is that large-scale datasets can be leveraged to learn high-level, platform and subject-agnostic signal representations. For EMG, such models could provide a unified backbone for many downstream tasks, building upon a robust, device-agnostic feature extraction, improving generalization across devices and subjects. However, progress in FMs of EMG has been limited by the scarcity of large sEMG corpora, and only with recent large-scale datasets (collected from thousands of users) [43] the exploration of generalized FMs of EMG became viable.
In this context, existing EMG FMs present significant limitations. Current models mostly focus on a single task (typically gesture classification) [10, 11]. Furthermore, the size of EMG FMs remains large, and no prior work demonstrated the deployment of an EMG FM on embedded hardware. This motivates the need for efficient FM architectures that meet the memory and computational constraints of embedded devices.
We address these limitations via the following contribution.
@@
  • Unlike previous EMG FMs, which are typically designed around a single downstream task or rely on large, domain-agnostic encoders, TinyMyo generalizes across heterogeneous sEMG datasets and sensor configurations. Our architecture follows the general Transformer design adopted in prior EMG works [10, 11], formulated in a unified EMG FM able to support a wide range of downstream tasks using the same pretrained backbone. We demonstrate the versatility of the pre-trained encoder through minimal task-specific head adaptations across diverse downstream tasks, including hand-gesture classification on the Ninapro DB5 [39], EPN-612 [6], UC Irvine [32] and Meta’s Generic Neuromotor Interface [26] datasets, hand-kinematic regression on the Ninapro DB8 [31], and speech production and discrimination on the Gaddy Silent Speech Dataset [16]. Our model achieves competitive performances across all downstream tasks, either surpassing SoA or guaranteeing a comparable accuracy with a sub-5M class model size.
  • We demonstrate for the first time the deployment of an EMG FM on an ultra-low-power microcontroller (GAP9), achieving an inference time of $0.785~\text{s}$, energy of $44.91~\text{mJ}$ and power envelope of $57.18~\text{mW}$.
  • We open source the model and the weights to enable reproducibility and further research by the community (https://github.com/pulp-bio/BioFoundation).
@@ 2 Related Works
Several efforts have been made to develop FMs for biosignals, and most of the works focused on electrocardiography (ECG), photoplethysmography (PPG), and electroencephalography (EEG) [37]. In contrast, research dedicated to EMG FMs remains limited.
@@
OTiS [50] is a 45 M parameters FM pre-trained on a broad collection of modalities, which demonstrated the feasibility of handling heterogeneous time-series data, including diverse input types, sampling rates, and channel configurations, within a unified pre-training framework. In the original study, EMG is used exclusively for multi-class muscular disease classification. Subsequent works by Cheng et al. [10, 11] evaluated the same model on hand-gesture classification using the EPN612 [6], NinaPro DB5 [39], NinaPro DB6 [38], and UCI EMG [32] datasets, demonstrating high accuracy across all datasets.
MOMENT [19] is a 350-M parameter FM pre-trained on a wide range of domains and temporal resolutions. While the original work did not consider EMG, recent works [11, 10] evaluated it for hand-gesture classification on the EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG datasets.
@@
PhysioWave [10] introduced a modality-specific FM for EMG, primarily evaluated on hand-gesture classification using the NinaPro DB5, EPN612, and UCI EMG. PhysioWave achieved a performance comparable to Moment and OTIS with a much smaller parameter count (5M for the smaller model).
WaveFormer [11] further refined this approach with a lightweight Transformer model (3.1 M parameters) tailored for sEMG-based gesture recognition. WaveFormer improved performance over PhysioWave on the EPN612, NinaPro DB5, NinaPro DB6, and UCI EMG datasets.
@@
Jiang et al. [24] introduced PhysioOmni, a multi-modal FM covering EEG, ECG, EOG and EMG, based on the same encoder as in LaBraM [25] (5.85.8M parameters per modality, considering its smallest-Base version). They adopt a decoupled multimodal tokenizer and masked signal modelling to ensure robustness under arbitrary missing modalities during inference. During pre-training, PhysioOmni relies on a fixed set of modalities (EEG, EOG, and EMG), which makes the architecture unsuitable for pre-training when one modality is missing. Downstream evaluations include sleep stage classification [2] and gait kinematic regression[9].
@@
In this context, existing EMG FMs are limited by the small number of explored downstream tasks, as well as by the lack of deployability on embedded platforms. TinyMyo is the first EMG FM to cover a broader range of downstream tasks with SoA performance, as will be extensively described in Section 4, while also being deployed on an ultra-low-power MCU.
@@ 3 Methods
Figure 1: Overview of the TinyMyo architecture. (a) Pre-training framework: the input signal is first tokenized using a channel-independent patching strategy with random masking, processed by eight bidirectional Transformer Encoder blocks, and reconstructed by a Lightweight Decoder. (b) Downstream architectures: after pre-training, the Decoder is removed and replaced with task-specific heads. For Hand-Gesture Classification and Kinematic Regression, a linear head is used to produce output classes or joint angles. For speech-related tasks, the EMG input is first downsampled through residual convolutional blocks and then passed through the pre-trained Transformer Encoder, followed by a linear head. The output is text for the Speech Recognition task. For the Speech Production task, a vocoder (HiFi-GAN) is added to produce audio from the predicted MFCC features. (c) Model Deployment: the architecture for the hand-gesture classification task is implemented on the GAP9 microcontroller.
@@ 3.1.1 Tokenization
We adopt a channel-independent patching strategy, preserving per-channel granularity at the embedding stage. Unlike EEG models that mix spatial information during tokenization via 2D convolutions spanning multiple channels [49] or project them into a unified latent space [13], we embed channels independently using a $1\times L$ kernel, and the subsequent transformer layers learn inter-channel dependencies via attention [12]. This design choice is grounded in the physiological nature of EMG recordings, where each channel corresponds to a distinct anatomical location.
Formally, given an input EMG signal $\mathbf{X}\in\mathbb{R}^{T\times C}$, with $T$ timesteps and $C$ channels, we apply a sliding window with patch length $L$ and stride $S$ independently to each channel. This yields a grid of patches $\mathbf{P}\in\mathbb{R}^{C\times N_{p}\times L}$, where $N_{p}=\lfloor(T-L)/S\rfloor+1$ is the number of temporal patches per channel.
@@
As input sequence for the Transformer, we flatten the channel and temporal dimensions, resulting in a total sequence length of $N=C\times N_{p}$. Each patch $\mathbf{P}_{c,i}$ ($c$ is the channel index and $i$ is the patch index) is mapped to a latent embedding of dimension $d_{e}$ via a shared learnable linear projection $\mathbf{W}_{\mathrm{proj}}\in\mathbb{R}^{d_{e}\times L}$. The resulting token sequence $\mathbf{E}\in\mathbb{R}^{N\times d_{e}}$ consists of embeddings $\mathbf{E}_{k}=\mathbf{W}_{\mathrm{proj}}\,\mathbf{P}_{c,i}^{\top}$ where the token index $k$ corresponds to the flattened arrangement of channel-time pairs $(c,i)$. We do not add learnable positional embeddings; instead, we inject positional information via Rotary Position Embeddings (RoPE) [47] within the attention blocks.
@@ 3.1.2 Architecture
The Encoder (see Fig. 1) comprises 8 layers with an embedding dimension of 192 and 3 attention heads. We utilize an MLP ratio of 4.0 (size 768), QKV bias (without QK norm), and a dropout rate of 0.1 for attention, projection, and drop-path. The lightweight Decoder shares the embedding dimension of 192.
We adopt a RoPE approach, which introduces position information by rotating query and key vectors in 2D planes instead of adding a position vector; such a method has already been adopted for biosignals [52, 10, 11].
During pre-training, the Encoder’s output is passed to a lightweight Decoder (3.9k parameters), consisting of a linear layer that outputs a sequence $\mathbf{\hat{P}}$, which is a reconstruction of the original patch sequence $\mathbf{P}$. Our approach draws inspiration from SimMIM [54], which demonstrates that a minimal reconstruction head (e.g., a single linear layer) is sufficient for effective self-supervised pre-training.
The use of a lightweight decoder is intentional: placing the full reconstruction burden on the encoder forces it to learn semantically rich and task-agnostic representations, rather than relying on a high-capacity decoder to compensate for missing information. This design strategy is supported by recent survey analyses in the masked image modeling literature: shallow decoders limit shortcut learning and encourage the formation of stronger latent representations within the encoder [59]. As a result, the learned features exhibit improved generalization across downstream tasks and recording conditions.
@@ 3.1.3 Pre-training objectives
We adopt a random-masking approach [54], where a random subset $\mathcal{M}$ of tokens is replaced by a learnable $\mathcal{[MASK]}$ token. Applying the masking over each patch enforces contextual inference over tens of milliseconds (physiologically meaningful burst segments), rather than trivial gap filling. We adopt a masking ratio of $50\%$ as in previous works [12, 49].
We employ the Smooth L1 Loss function as reconstruction objective between masked and original patches, as in [10]. We define the following loss components:
@@
All experiments are implemented in Python 3.10 using PyTorch Lightning and Hydra for modularity, configurability, and reproducibility. The proposed foundation model is trained on the CSCS Alps HPC infrastructure using NVIDIA GH200 GPUs, employing a single node with 4x NVIDIA GH200 GPUs in Distributed Data Parallel (DDP) mode. We train using the AdamW optimizer ($\beta_{1}=0.9,\beta_{2}=0.98$, weight decay 0.01) with a global batch size of 512, utilizing gradient accumulation over 8 batches and a clipping threshold of 3. The learning rate follows a cosine schedule for 50 epochs (10 warm-up epochs), peaking at $1\times 10^{-4}$ (minimum $1\times 10^{-6}$). The model processes sequences of 1000 timesteps with a masking ratio of 0.50 .